# Dense Backpropagation Improves Training for Sparse Mixture-of-Experts

### v2

@@ 4.3 Ablations
(a) Hidden Dimension = 512
(b) Hidden Dimension = 1024
(c) Hidden Dimension = 2048
Figure 6: Ablating the hidden dimension. We perform an ablation study across varying hidden dimensions for both the Default MoE and a Top-K baseline approach. We plot the perplexity across 3 hidden dimensions: 512, 1024 and 2048. Notably, the Default MoE consistently outperforms the Top-K baseline across all model sizes, demonstrating lower perplexity at every scale.
Default MoE Remains Superior As We Increase the Model Size.  Figure 6 compares the two methods as we increase the size of the model from $557M$ total parameters at a hidden dimension of $512$, to $7.33B$ total parameters when the hidden dimension is $2048$. Our method outperforms the baseline across all model sizes. All other results in the paper use a hidden dimension of $1024$.
MoE Architecture Ablations. We apply Default MoE with different expert configurations corresponding to various levels of sparsity, using NcK to refer to a model with $N$ total experts and $K$ active experts. When increasing the number of experts past $8$, we fine-grain experts (Dai et al., 2024) so that the total number of model parameters remains constant. Specifically, the intermediate size is now $704$ rather than $2816$ because we have 4×4$\times$4 × the number of experts). In Figure 8, we train both a standard Top-K model and Default MoE with the configurations 8c1, 8c2, 32c1, 32c2, and 32c4, which correspond to sparsity factors of $1/8$, $1/4$, $1/32$, $1/16$ and $1/8$ respectively.
@@
Default MoE outperforms a standard TopK MoE at all sparsity configurations. Our improvement is more pronounced for the lowest sparsity (8c2) but still significant for higher sparsity MoEs such as 32c1. Notably, Default MoE takes longer to ”warm up” for sparser models; for example, in Figure 8 the 32c1 Default MoE rapidly catches up and surpasses the TopK MoE around the 10 billion token mark, while the 32c2 MoE is clearly ahead of TopK even at 2 billion tokens.
While we use $\beta=0.9$ for the 8c1 and 8c2 MoEs in Figure 8, $\beta$ requires more careful tuning for the 32 expert MoEs. For $N=32$, we use $\beta=0.65$, $\beta=0.95$, and $\beta=0.999$ for $K=1$, $K=2$, and $K=4$ respectively. In other words, sparser MoEs require a lower $\beta$. We believe this is due to each expert receiving less tokens at each step, which leads to a sparser ”history” for each default vector. As a result, the default vector’s estimate for the average expert output improves when assigning higher weight to the current batch.
Figure 7: Comparison of Default MoE and Top-K Across Different Configurations. We compare Default MoE and Top-K across five configurations: 8c1, 8c2, 32c1, 32c2, and 32c4. The results show that Default MoE outperforms Top-K in all MoE configurations. Both 8c1 and 8c2 use $\beta=0.9$; however, 32c1, 32c2, and 32c4 require more careful tuning of $\beta$. We use $\beta=0.65$, $\beta=0.95$, and $\beta=0.999$ for 32c1, 32c2, and 32c4, respectively. We detail our choices of $\beta$ in Section A.1.
Figure 8: Learning Rate Sweep for Top-K and Default MoE. We perform a learning rate sweep for the Top-K approach to determine the largest stable learning rate, which is then used for Default MoE. The results show that while Top-K struggles with higher learning rates—evidenced by poor performance at $9\times 10^{-4}$—Default MoE remains stable and can effectively utilize larger learning rates. This is demonstrated by Default MoE achieving comparable performance at $9\times 10^{-4}$ to Top-K at $7\times 10^{-4}$, highlighting its greater training stability.
@@
Tuning the Learning Rate. We want to ensure that we are comparing against a tuned baseline, so we tune the learning rates in Figure 8. We use the learning rate that achieves the best performance for the TopK baseline, $7\times 10^{-4}$, for our main results in Figure 4. However, this is actually not the best learning rate for the Default MoE; the larger learning rate of $9\times 10^{-4}$ is slightly better. By contrast, $9\times 10^{-4}$ is much too large for the baseline. This indicates that our Default MoE is more stable to train, likely because we are updating the entire router. If we train the baseline with such a large learning rate, we see that a single iteration with a very imbalanced load will lead to a large loss spike; this iteration is very noticeable because, since we train dropless MoEs, it is also slower. We never observe this for the Default MoE. It is unsurprising that our Default MoE’s best learning rates may be different from the baseline, but importantly, our method outperforms the baseline at all learning rates we consider.
@@
We present additional ablations regarding the EMA, including the hyperparameter $\beta$, the EMA initialization, and passing the EMA value forward, in Section A.1.
@@ 4.4 Efficiency
Table 3: Throughput Comparison. We compare the throughput between our method and the TopK baseline for different model sizes. Throughput is measured in tokens per second (sequence length 2048) on a single GPU.
@@
Our Method Does Not Significantly Reduce Throughput. In Table 3 we report the throughput in samples per second and of TopK and our method while training the standard 8-expert MoE on a single GPU node. Our method is quite lightweight, because we just need to update an EMA in the forward pass and then update the router gradients in the backward pass. However, for the 2B model we train, this is still nearly a $2\%$ overhead; less than the $9\%$ speedup we observe over the TopK baseline in Figure 4, but still non-zero. We find that the overhead of our method decreases as we increase the hidden size. The 2B model we train has $1024$ hidden size, $2048$ hidden size is an 2B model, etc. As the hidden size increases, the proportion of time spent in the MLP matmuls increases, and the overhead of our model is no longer significant compared to the total MoE layer runtime. For even just an 8B model, our overhead is near-zero; and of course, this is almost 100×100$\times$100 × smaller than production MoEs (DeepSeek-AI Team, 2024b). We are confident that our lightweight EMA method does not reduce throughput at scale.
Our Method Does Not Significantly Increase the Memory Footprint. The EMA buffers are not updated during the backward pass, so we don’t need to store any additional activations. The only additional memory required by our method is the EMA buffers themselves. For each expert in each layer, we store a buffer of the same size as the model’s hidden dimension. The total number of parameters in an expert is $hidden\_size\times intermediate\_size$; for our MoE this is $1024\times 2816$. And we are just increasing this by $1024$, which is a $1/2816=0.03\%$ increase in the number of parameters. The MoE parameters are not the entirety of the model (they are about $3/4$) so the additional memory footprint of our method is negligible.
@@ 5 Discussion
We propose a training method to improve language modeling performance for MoEs. By approximating the signal of a dense mixture-of-experts layer, the MoE router is able to receive information from all experts for each token, regardless of how many were activated. This approximated dense signal allows our Default MoE to outperform the standard Top-K baseline in a variety of settings. Moreover, the Default MoE requires minimal added computational overhead which decreases as the model size grows larger, showcasing the potential of our method for large-scale MoE pretraining.
6 Acknowledgements
We would like to thank the Capital One AGI Foundations Team and in particular Andrei Mircea, Stephen Rawls, and Sambit Sahu for helpful discussions and support for the paper. This material is based upon work partially supported by the NSF Grant No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Appendix
A.1 Ablating EMA design choices
Tuning $\beta$. The lone hyperparameter introduced by our method is the $\beta$ parameter of the EMA. We vary the $\beta$ used in the Default MoE in Figure 9 for an 8 expert MoE and find that both $\beta=0.9$ and $\beta=0.999$ perform equally well. Intuitively, $\beta$ parametrizes how fast our EMA adjusts its model of the sample mean as we train the model. In principle the optimal value of $\beta$ might depend on the learning rate, which defines how fast the model is changing, the batch size, which controls how many tokens are actually incorporated into the EMA, and even the data distribution. For example, we need to tune $\beta$ more carefully for a 32 expert MoE. In Figure 10 we sweep across multiple values of $\beta$ for a 32c1, 32c2, and 32c4 Default MoE.
@@ A.1 Ablating EMA design choices
Initializing the EMA. We experiment with initializing default vector EMA in two ways: zero initialization and random (Gaussian) initialization. Zero initialization leads to the default vector starting off with zero signal in the earliest training steps. Random initialization, on the other hand, immediately provides some signal to fill in missing expert outputs but this signal starts as pure noise. Given these tradeoffs, we try both approaches and outline our results in Figure 11. Initializing our EMA with zeroes demonstrates slight improvement compared to random initialization. We believe this is due to the adverse effects from providing a random signal to the router early in training. However, the impact is minor, even without the use of any explicit bias correction term in our EMA.
@@
Passing the Default Vector Forward. Our goal is to provide a default expert activation for unactivated experts so that the router can receive a gradient for all experts. We do this by passing the default vector forward through the network, and the router gradient is automatically computed. However, we can in principle do this without passing the default vector forward, and just manually writing the gradient update for the backward pass. One reason why we pass the default vector forward is because our error in estimating the true dense gradient, as written in Eq. 9, is scaled by the loss. Therefore if the default vector activations actually improve the model’s output, our error is smaller. We validate this in Figure 12, where we find that passing the default vector forward does improve over only updating the router gradients in the backward pass with the default vector.
A.2 Note on References

### v3

@@ 4.3 Ablations
Tuning $\beta$. We initially tuned $\beta$ for every configuration independently, and the results of this can be found in Section˜A.4. We found that more sparse configurations, such as 32c1, required a lower $\beta=0.65$ whereas a much higher beta of $\beta=0.999$ was necessary for 32c4 to achieve its best performance. However, we can actually automatically account for the impact of sparsity and granularity on the default vector update by weighting the updates to the default vector by the router logits. When we do this, different values of $\beta$, that would otherwise be suboptimal, all converge to the same good performance. Therefore, in our final method, the value of $\beta$ does not seem to matter as shown in  Figure˜8.
@@
MoE Architecture Ablations. In Figure˜2(a), we train both a standard TopK model and Default MoE with the configurations 8c1, 8c2, 32c1, 32c2, and 32c4, which correspond to sparsity factors of $1/8$, $1/4$, $1/32$, $1/16$ and $1/8$ respectively. Default MoE outperforms a standard TopK MoE at all sparsity configurations. Our improvement is more pronounced for the lowest sparsity (8c2) but still significant for higher sparsity MoEs such as 32c1. Notably, Default MoE takes longer to "warm up" for sparser models; for example, we notice in Figure˜13 that the 32c1 Default MoE rapidly catches up and surpasses the TopK MoE around the 10 billion token mark, while the 32c2 MoE is clearly ahead of TopK even at 2 billion tokens.
@@
Tuning the Learning Rate. We want to ensure that we are comparing against a tuned baseline, so we tune the learning rates in Figure˜2(b). We use the learning rate that achieves the best performance for the TopK baseline, $7\times 10^{-4}$, for our results in Figure˜12. We observe that while it performs better than Default MoE earlier in training (across our learning rate sweep experiments), the results in  Figure˜12 show that even in this case, Default MoE eventually surpasses TopK. Moreover, we note that the best learning rate for the Default MoE is the larger learning rate of $9\times 10^{-4}$. By contrast, $9\times 10^{-4}$ is much too large for the baseline. This indicates that our Default MoE is more stable to train, likely because we are updating the entire router. If we train the baseline with such a large learning rate, we see that a single iteration with a very imbalanced load will lead to a large loss spike; this iteration is very noticeable because, since we train dropless MoEs, it is also slower. We never observe this for the Default MoE. It is unsurprising that our Default MoE’s best learning rates may be different from the baseline, but importantly, our method outperforms the baseline at all learning rates we consider.
Default MoE Remains Superior As We Increase the Model Size.  Figure˜2(c) compares the two methods as we increase the size of the model from $557M$ total parameters at a hidden dimension of $512$, to $7.33B$ total parameters when the hidden dimension is $2048$. Our method outperforms the baseline across all model sizes. All other results in the paper use a hidden dimension of $1024$.
(a) MoE Configurations
(b) Learning Rate Sweep
(c) Ablation over model parameters
(d) Sparsity Factor Ablation
Figure 2: Comprehensive Ablations. We report the perplexity after training for 10B tokens for Default MoE and Top-K while varying: (a) MoE configuration across five MoE setups. (b) Learning rate while keeping the MoE configuration at 8c1. (c) Total model size while keeping the MoE configuration at 8c1. (d) Sparsity, where the final 1/64 sparsity is for a 7B model, therefore the perplexity is lower than the other sparsity factors that are for 2B models.
@@
We present additional ablations regarding the EMA, including the hyperparameter $\beta$, the EMA initialization, and passing the EMA value forward, in Section˜A.4.
4.4 Comparison to Other Routing Methods
We consider three alternatives to TopK routing for comparison: Sparesemixer (Liu et al., 2023, 2024), ReMoE (Wang et al., 2025), and Loss-Free Balancing (Wang et al., 2024).
Default MoE Beats Sparsemixer.  Liu et al. (2023) proposes Sparsemixer, which estimates the true router gradient without straight-through.  Liu et al. (2024) note that Sparsemixer lags behind TopK (which our method always outperforms) for the first $0.5T$ tokens, likely due to the noise that Sparsemixer adds. After 24B tokens, Sparsemixer obtains a validation perplexity of $19.77$, worse than both TopK and our method. We present the training curve in Figure˜5.
ReMoE Fails to Converge. We cannot compare our zero-shot benchmark scores to ReMoE because they train on The Pile, which is unavailable. We train ReMoEs on FineWeb-Edu. ReMoE reports that the first stage of their method is expected to run slowly while their method starts to work, as it is essentially acting as a dense model for the first few hundred steps. Following that, they report that ReMoE exhibits a significant loss spike around 2 billion tokens, after which it recovers and starts to outperform the baseline. In our setting, we indeed were able to replicate the first two results, where ReMoE runs very slowly while acting as a dense model, but the sparsity coefficients start to converge, and the loss spikes at 2 billion tokens. However, the loss never recovered. We invested over 1000 GPU hours into trying to get the ReMoE model to train, but did not have any success.
Loss-Free Balancing Underperforms. We first sweep the bias update coefficient, which is the hyperparameter in the loss-free balancing method used in Wang et al. (2024), and find that $0.0001$ provides the best results. However, after training the 8c1 MoE for 10B tokens, Loss-Free Balancing obtains a validation perplexity of $24.22$, whereas our Default MoE obtains $23.48$.
Why do Prior Methods Underperform? Prior work claims to improve upon standard dMoEs, but there is one difference in our experimental setup: we use globally reduced load balancing loss (Qiu et al., 2025) in all experiments. Corroborating their conclusions, we find that globally reducing the auxiliary loss significantly improves the load balancing, and subsequently the performance, of the baseline, therefore making the improvement of prior methods less noticeable. This is because the “free lunch” of slightly improving routing is more or less eaten by the global auxiliary loss reduction, which is trivial to implement and uses little compute, and prior work frequently targets this same free lunch as their source of improvement. Qiu et al. (2025) directly evaluate Loss-Free Balancing and report that it underperforms a proper globally reduced auxiliary loss, supporting our results.
4.5 Empirical Analysis
In Appendix˜B we validate that the load balance, expert coactivation, and domain specialization of the trained MoEs do not deteriorate when we implement our DefaultMoE. Our results corroborate those of  Qiu et al. (2025), that even a baseline model trained with globally reduced auxiliary loss will have good load balance and domain specialization.
Figure 3: Analysis. Left: Router Entropy after training for 10B tokens. Middle, Right: Heatmap of the cosine similarities between default vectors for the 8c1 DefaultMoE at layers 0 and 16.
In Figure˜3(a) we observe that the router entropy is highest for the early layers and consistently goes down, with the final layer’s router entropy being very low. Intuitively, the router shouldn’t quite know where to send a token at the very start of the model, but the routing decisions should become more and more fixed as we proceed through the model.
Similarity of Default Vectors. In Figure˜3(a) we plot heatmaps of the cosine similarities between default vectors for experts in different layers of the 8c1 DefaultMoE. The default vectors at layer 0 are not very similar, indicating that the default vectors are learning something nontrivial. However, at layer 16 the default vectors are actually very similar. This follows the trend of the decline in router entropy of Figure˜3; if the router is very confident where a token should go, all the default vectors may learn similarly trivial things, but if the router has high entropy, the default vectors should learn faithful nontrivial approximations of the expert activations.
Figure 4: Dense Router Gradient Similarity. We plot similarity to the dense router gradient (K=8) for our 8ck Top-K and DefaultMoEs pretrained on 160B tokens. The DefaultMoE’s router gradient is much more similar to the dense router gradient.
Similarity to Dense Gradient. Our entire design of DefaultMoE is based on the premise that we can better approximate the dense gradient corresponding to simultaneous activation of all the 8 experts. In Figure˜4, we validate that our DefaultMoE’s router gradient is more similar to the dense gradient than the TopK MoE’s router gradient is. We take the 8c1 MoE pretrained on 160B tokens and compute the dense router gradient, by activating all the 8 experts. We then vary K between 1 and 7 (the full range can be found in Appendix˜B) and report the similarity of the sparse router gradient to the dense gradient, for both DefaultMoE and TopK MoE. At K=1, DefaultMoE is clearly much more similar to the dense gradient for the first few layers, while TopK is not close at all. This is because the routing distribution has high entropy at the early layers, so in DefaultMoE, the early layers will make significant use of the default vectors. This increases the gradient flow to the router which is reflected in the closer approximation of the dense gradient. Whereas even if the router is allocating near-equal mass on all experts and TopK MoE only selects one, its router gradient will be quite dissimilar from the dense gradient. Even at K=7, TopK MoE is still not quite approximating the dense gradient, indicating importance of the missing information from the least activated expert.
@@ 4.6 Efficiency
Our Method Does Not Significantly Reduce Throughput. When we train the 7.33B 8-expert MoE, the per-node GPU throughput is 1,393 tokens per second for TopKMoE and 1,391 tokens per second for DefaultMoE, with a variation of about 1-2 tokens per second on each of these numbers; this is a throughput decrease of just $0.18\%$ which is just variance. For larger models, more of the time will be spent in matmuls and the overhead of our method will be insignificant.
Our Method Does Not Significantly Increase the Memory Footprint. The only additional memory required by our method is the EMA buffers themselves. For each expert in each layer, we store a buffer of the same size as the model’s hidden dimension. The total number of parameters in an expert is $hidden\_size\times intermediate\_size$; for our MoE this is $1024\times 2816$. We increase this by $1024$, which is a negligible $1/2816=0.03\%$ increase in the number of MoE parameters.
Table 3: Throughput Comparison. We compare the throughput between DefaultMoE and TopKMoE for different model sizes, measured in tokens per second (sequence length 2048) on 1 GPU.
@@ 5 Discussion
We have validated that DefaultMoEs beat tuned TopKMoEs across multiple settings, but why? Our explanation is as follows. Our default vectors are serving as nontrivial (Figure˜3) approximations of their respective expert activations. Our gradient error should be corrected, which is validated by DefaultMoE’s router gradient being closer to the dense router gradient than that of TopKMoE (Figure˜4). Less gradient error means that DefaultMoE can tolerate a higher learning rate than TopKMoE (Figure˜2), or simply converge faster at TopKMoE’s best learning rate (Figure˜12) and score higher on benchmarks (Table˜2). This improvement is efficient (Table˜3), and does not require additional hyperparameter tuning (Figure˜8). We hope that our work will help the community train better MoEs.
@@ Appendix A Appendix
A.1 Limitations
Throughout our work we only train dropless MoEs. These are ideal for the academic research setting, because they train efficiently in data-parallel and so do not require cumbersome implementations of model or pipeline parallelism. However, dropless MoEs may not be well-suited for industry, because reshaping the tensors to provide consistent shapes to the MoE kernels incurs many host-device syncs. This is a limitation that we share with prior work, but an important future direction will be extending our results to token-dropping MoEs.
We report results on a range of standard language modeling benchmarks, but we have not invested enough compute to see performance on LLM benchmarks such as MMLU-Pro, and our data is not sufficiently domain-specific to report scores on coding benchmarks (HumanEval, MBPP) or math (GSM8k). In the same vein, our models have not been trained to reason. We focus heavily on pretraining in this paper, but an important future direction would be seeing whether we can finetune MoEs with our method.
A.2 Experimental Setup Details
Architecture. We use SwiGLU [Shazeer, 2020] MLPs following Llama [Touvron et al., 2023], $16$ attention heads with dimension of $64$, LayerNorm [Ba et al., 2016] and RoPE [Su et al., 2023].
Hyperparameters. We use the initialization from  Wang et al. [2022] for residual branch merge layers and the initialization from Nguyen and Salazar [2019] for all other layers. We use the AdamW optimizer [Loshchilov and Hutter, 2019]. We use a sequence length of $2048$ and a global batch size of $1024$, resulting in a global token batch size of $2^{21}$. We use a standard cosine decay schedule.
MoE-Specific Training Details. We set the auxiliary loss [Fedus et al., 2022] to $0.01$. We do not use z-loss or jitter, because we find that at this scale they do not improve training for the baseline model. Following DeepSeek-AI Team [2024b], we set the first layer to be dense. Following Liu et al. [2024], we compute the aux loss across nodes. We train dropless MoEs.
Implementation. We train with the gpt-neox library [Andonian et al., 2023] integrated with Megablocks [Gale et al., 2022] and augmented with Triton kernels from [Hsu et al., 2024].
A.3 More details on other Routing Mechanism Comparisons
We see in Figure˜5 that our method significantly outperforms SparseMixer for at least the first $10B$ tokens.
A.4 Ablating EMA design choices
Tuning $\beta$. The lone hyperparameter introduced by our method is the $\beta$ parameter of the EMA. While we use $\beta=0.9$ for the 8c1 and 8c2 MoEs in Figure˜4, $\beta$ requires more careful tuning for the 32 expert MoEs. For $N=32$, we use $\beta=0.65$, $\beta=0.95$, and $\beta=0.999$ for $K=1$, $K=2$, and $K=4$ respectively. In other words, sparser MoEs require a lower $\beta$. We believe this is due to each expert receiving less tokens at each step, which leads to a sparser "history" for each default vector. As a result, the default vector’s estimate for the average expert output improves when assigning higher weight to the current batch.
We vary the $\beta$ used in the Default MoE in Figure˜6 for an 8 expert MoE and find that both $\beta=0.9$ and $\beta=0.999$ perform equally well. Intuitively, $\beta$ parametrizes how fast our EMA adjusts its model of the sample mean as we train the model. In principle the optimal value of $\beta$ might depend on the learning rate, which defines how fast the model is changing, the batch size, which controls how many tokens are actually incorporated into the EMA, and even the data distribution. For example, we need to tune $\beta$ more carefully for a 32 expert MoE. In Figure˜7 we sweep across multiple values of $\beta$ for a 32c1, 32c2, and 32c4 Default MoE.
@@ A.4 Ablating EMA design choices
Figure 8: We compare exponential moving average (EMA) updates to the default vector with and without scoring. In the scored version, each expert’s contribution is weighted by its router probability before being added to the EMA update. Without scoring, the choice of $\beta$ (e.g., 0.9 vs. 0.999) significantly affects performance. With scoring, this sensitivity disappears—both scored variants converge to the same strong final performance. The unscored $\beta=0.999$ curve (orange) is nearly hidden behind the scored curves.
Weighted Updates Remove Sensitivity to EMA Decay Rate ($\beta$). We explore different values of the EMA decay coefficient $\beta$ used to update the default vector. Naively, performance is sensitive to $\beta$. However, we find that if we accumulate activations in the the default vector weighted by their router score performance becomes robust to the choice of $\beta$. In fact, multiple $\beta$ values converge to the a similar final performance as shown in  Figure˜8.
@@
Initializing the EMA. We experiment with initializing default vector EMA in two ways: zero initialization and random (Gaussian) initialization. Zero initialization leads to the default vector starting off with zero signal in the earliest training steps. Random initialization, on the other hand, immediately provides some signal to fill in missing expert outputs but this signal starts as pure noise. Given these tradeoffs, we try both approaches and outline our results in Figure˜9. Initializing our EMA with zeroes demonstrates slight improvement compared to random initialization. We believe this is due to the adverse effects from providing a random signal to the router early in training. However, the impact is minor, even without the use of any explicit bias correction term in our EMA.
@@
Passing the Default Vector Forward. Our goal is to provide a default expert activation for unactivated experts so that the router can receive a gradient for all experts. We do this by passing the default vector forward through the network, and the router gradient is automatically computed. However, we can in principle do this without passing the default vector forward, and just manually writing the gradient update for the backward pass. One reason why we pass the default vector forward is because our error in estimating the true dense gradient, as written in Eq. 9, is scaled by the loss. Therefore if the default vector activations actually improve the model’s output, our error is smaller. We validate this in Figure˜10, where we find that passing the default vector forward does improve over only updating the router gradients in the backward pass with the default vector.
Figure 13: Comparison of MoE Configurations. We compare Default MoE and Top-K across five configurations: 8c1, 8c2, 32c1, 32c2, and 32c4. The results show that Default MoE outperforms Top-K in all MoE configurations. Both 8c1 and 8c2 use $\beta=0.9$; however, 32c1, 32c2, and 32c4 require more careful tuning of $\beta$. We use $\beta=0.65$, $\beta=0.95$, and $\beta=0.999$ for 32c1, 32c2, and 32c4, respectively. We detail our choices of $\beta$ in Section˜A.4.
Appendix B Detailed Empirical Analysis
B.1 Domain Specialization
We plot the domain specialization for 8c1 TopKMoE (Figure˜14,Figure˜15,Figure˜16), 8c1 DefaultMoE (Figure˜20,Figure˜21, Figure˜22), 32c4 TopKMoE (Figure˜17,Figure˜18,Figure˜19) and 32c4 DefaultMoE (Figure˜23,Figure˜24,Figure˜25). All models are pretrained on FineWeb-Edu for 160B tokens.
Figure 14: Domain Specialization for 8c1 TopKMoE at Layer 0.
Figure 15: Domain Specialization for 8c1 TopKMoE at Layer 8.
Figure 16: Domain Specialization for 8c1 TopKMoE at Layer 16.
Figure 17: Domain Specialization for 32c4 TopKMoE at Layer 0.
Figure 18: Domain Specialization for 32c4 TopKMoE at Layer 8.
Figure 19: Domain Specialization for 32c4 TopKMoE at Layer 16.
Figure 20: Domain Specialization for 8c1 DefaultMoE at Layer 0.
Figure 21: Domain Specialization for 8c1 DefaultMoE at Layer 8.
Figure 22: Domain Specialization for 8c1 DefaultMoE at Layer 16.
Figure 23: Domain Specialization for 32c4 DefaultMoE at Layer 0.
Figure 24: Domain Specialization for 32c4 DefaultMoE at Layer 8.
Figure 25: Domain Specialization for 32c4 DefaultMoE at Layer 16.
B.2 Load Balancing
We plot the load balancing for 8c1 TopKMoE (Figure˜26), 32c4 TopKMoE (Figure˜27), 8c1 DefaultMoE (Figure˜28) and 32c4 DefaultMoE (Figure˜29). All models are pretrained on FineWeb-Edu for 160B tokens.
Figure 26: Routing Frequency for 8c1 TopKMoE.
Figure 27: Routing Frequency for 32c4 TopKMoE.
Figure 28: Routing Frequency for 8c1 DefaultMoE.
Figure 29: Routing Frequency for 32c4 DefaultMoE.
B.3 Expert Coactivation
We plot the expert coactivation for 32c4 DefaultMoE at Layer 0 (Figure˜30) and Layer 16 (Figure˜31). All models are pretrained on FineWeb-Edu for 160B tokens.
Figure 30: Expert Coactivation for 32c4 DefaultMoE at Layer 0.
Figure 31: Expert Coactivation for 32c4 DefaultMoE at Layer 16.
B.4 Dense Gradient Similarity
We plot the similarity to the dense gradient for 8c1 TopKMoE and DefaultMoE at K=1 (Figure˜32), K=2 (Figure˜33), K=3 (Figure˜34), K=4 (Figure˜35), K=5 (Figure˜36), K=6 (Figure˜37) and K=7 (Figure˜38). All models are pretrained on FineWeb-Edu for 160B tokens.
Figure 32: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=1.
Figure 33: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=2.
Figure 34: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=3.
Figure 35: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=4.
Figure 36: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=5.
Figure 37: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=6.
Figure 38: Similarity to Dense Gradient for 8c1 TopKMoE and DefaultMoE at K=7.
B.5 Default Vector Similarities
We plot the cosine similarity of default vectors in 8c1 DefaultMoE (Figure˜39, Figure˜40, Figure˜41) and 32c4 DefaultMoE (Figure˜42, Figure˜43, Figure˜44). All models are pretrained on FineWeb-Edu for 160B tokens.
Figure 39: Similarities Between Default Vectors for 8c1 DefaultMoE at Layer 0.
Figure 40: Similarities Between Default Vectors for 8c1 DefaultMoE at Layer 8.
Figure 41: Similarities Between Default Vectors for 8c1 DefaultMoE at Layer 16.
Figure 42: Similarities Between Default Vectors for 32c4 DefaultMoE at Layer 0.
Figure 43: Similarities Between Default Vectors for 32c4 DefaultMoE at Layer 8.
Figure 44: Similarities Between Default Vectors for 32c4 DefaultMoE at Layer 16.
B.6 Note on References